OpenStudy (anonymous):

What are the exact solutions of x^2+5x+1-0?

OpenStudy (anonymous):

x²-5x-1 = 0 Use quadratic formula: a = 1, b = -5, c = -1 x = [-b±√(b²-4ac)]/(2a) x = {-(-5)±√[(-5)²-4(1)(-1)]}/[2(1)] x = {5±√[(25)+(4)]}/2 x = [5±√(29)]/2 Solutions: x = [5±√(29)]/2

OpenStudy (anonymous):

b and c are 5 and1repectively